A business firms sells a good at the price of Rs 450.The firm has decided to reduce the
price of good to Rs 350.Consequently, the quantity demanded for the good rose from
25,000 units to 35,000 units. Calculate the price elasticity of demand.
1
Expert's answer
2020-12-25T15:26:16-0500

"\bold {ANSWER}"


The price elasticity of demand, "\eta_{d} = -1.40"


"\bold {SOLUTION}"

We are given:

"P_{0} = \$450" , "Q_{0} = 25,000 \space units"

"P_{1} = \$350" , "Q_{1} = 35,000 \space units"


"\eta_{d} = \dfrac {\% \space ∆ \space in \space Q_{d}}{\% \space ∆ \space Price}"


"\% \space ∆ \space in \space Qd =" "\dfrac {35,000 - 25,000}{25,000}×100\%"


"= \dfrac {10,000}{25,000}×100\%"


"= 40\%"



"\% \space ∆ \space in \space Price = \dfrac {\$350 - \$450}{\$350}×100\%"


"= \dfrac {-\$100}{\$350}×100\%"


"=- 28.571429\%"



"\therefore \space \eta_{d} = \dfrac {40\%}{-28.571429\%}"


"= -1.39999999988"


"= \bold {-1.40}"


Thus, the price elasticity of demand is -1.40